# A Multi-scale Information Integration Framework for Infrared and Visible Image Fusion

### v1

@@ Abstract
Infrared and visible image fusion aims at generating a fused image containing the intensity and detail information of source images, and the key issue is effectively measuring and integrating the complementary information of multi-modality images from the same scene. Existing methods mostly adopt a simple weight in the loss function to decide the information retention of each modality rather than adaptively measuring complementary information for different image pairs. In this study, we propose a multi-scale dual attention (MDA) framework for infrared and visible image fusion, which is designed to measure and integrate complementary information in both structure and loss function at the image and patch level. In our method, the residual downsample block decomposes source images into three scales first. Then, dual attention fusion block integrates complementary information and generates a spatial and channel attention map at each scale for feature fusion. Finally, the output image is reconstructed by the residual reconstruction block. Loss function consists of image-level, feature-level and patch-level three parts, of which the calculation of the image-level and patch-level two parts are based on the weights generated by the complementary information measurement. Indeed, to constrain the pixel intensity distribution between the output and infrared image, a style loss is added. Our fusion results perform robust and informative across different scenarios. Qualitative and quantitative results on two datasets illustrate that our method is able to preserve both thermal radiation and detailed information from two modalities and achieve comparable results compared with the other state-of-the-art methods. Ablation experiments show the effectiveness of our information integration architecture and adaptively measure complementary information retention in the loss function. Our code is available at https://github.com/SSyangguang/MDA.
Index Terms:
@@
infrared and visible image fusion, dual attention, adaptive weight generation.
I Introduction
@@
Due to the limitations of different physical sensors, images from one single sensor is hard to comprehensively describe the scene information. Image fusion is one kind of multimodal information fusion technique, which aims at integrating images from different modalities into a single more informative image at the pixel level [1, 2]. Infrared and visible image fusion is a branch of the image fusion techniques to generate an image containing both thermal and detail information, in which the infrared images can easily distinguish targets with high temperature from the background [3]. However, infrared images typically have low spatial resolution and poor details. As complementary information to the infrared images, the visible images contain rich textures but are sensitive to unfavorable illumination conditions like nighttime and bad weather. The fusion results of infrared and visible images can depict the scene from aspects of thermal radiations and the human visual system simultaneously, which is practical in many fields, such as recognition, surveillance and object detection [4, 5, 6].
@@
According to the adopted theories [7], the traditional image fusion method can be divided into several categories, namely multi-scale transform-, sparse representation-, subspace-, saliency-bases methods, hybrid models, and other models. However, the fusion performance of these methods is highly related to their handcrafted features and fusion schemes [8]. In recent years, deep learning-based methods have been successfully applied to image fusion with their strong feature extraction ability and adaptive fusion strategy, which overcome the drawbacks of traditional methods to a certain extent. The overall deep learning-based image fusion pipeline can be processed implicitly by an end-to-end convolutional neural network under the well-designed network architecture and loss function, which is able to characterize complex relationships between input image pairs and the fusion image[9]. As suggested by [10], the deep learning-based image fusion methods can be divided into autoencoder (AE)-based methods, convolutional neural network (CNN)-based methods and generative adversarial network (GAN)-based methods. Among them, GAN-based methods rely on an adversarial model between the generator and discriminator to implicitly fulfill feature extraction, feature fusion and image reconstruction, such as FusionGAN [11]. The AE-based and CNN-based methods generally use an autoencoder or fully convolutional network to realize the feature extraction and reconstruction, while the fusion strategy is conventional fusion rules or concatenation on channel dimension, such as DenseFuse [12] and PMGI [13].
@@
Infrared and visible images capture radiation and illumination information respectively, and typically pixel intensity is used to constrain thermal radiation information retention while gradient variation is used to constrain texture detail information [14]. Nevertheless, the complementary information varies greatly in different scenes. For instance, some infrared images achieve better detail at night, which makes the above information measurement method not accurate enough [15]. Existing methods prefer preserving pixel intensity and detail information via the mean square error and structural similarity separately, and most of them simply add residual or dense connection in the network architecture to capture texture details, failing to measure the complementary information among different modalities, which is a key issue in image fusion [16]. Additionally, no ground truth in the infrared and visible image fusion task makes manually designed information retention methods unreasonable. Therefore, determining the degree of vital information retention based on each modality’s own attributes and effectively integrating complementary information makes infrared and visible image fusion more reasonable. The existing methods barely applied complementary information measurement neither in calculating loss function nor feature fusion module.
@@
[17] showed that multi-scale analysis is similar to the human visual characteristics, and the fused image with such property is capable of capturing contextual information from the high-level representations and detailed information from the low-level representations [18]. In this paper, we propose a multi-scale dual attention (MDA) framework applying complementary information both in loss function and network architecture to effectively measure and integrate vital information at three scales. The given images are first decomposed into three scales via the residual downsample block. High-resolution features are more beneficial for measuring and preserving detailed information, while the pixel intensity distribution can be measured accurately at low resolution. Differing from those directly aggregate features by addition or concatenation, we introduce channel and spatial attention mechanisms to exploit ’what’ and ’where’ is informative for the given infrared-visible image pairs [19, 20], and then fuse features extracted by the residual downsample block based on the attention maps. Dual attention fusion block merges feature from the same and adjacent scale of different modalities along the spatial and channel dimension, each of them generating an attention map for feature aggregation, which is considered the complementary information aggregation process. Fused features are upsampled to the same size as the input image and reconstructed by the residual reconstruction block.
@@
As an auxiliary to measuring the multi-scale complementary information, the pixel item in the loss function consists of intensity and detail two parts with different adaptive weight coefficients obtained by the complementary information measurement, based on the full resolution inputs and local patch. Additionally, we add a style loss to constrain pixel intensity distribution differences between the fusion image and the infrared image. The main contributions in this work are as follows:
@@
    We design a multi-scale dual attention (MDA) framework for infrared and visible image fusion that extracts features from different modalities across multiple spatial scales, both utilizing pixel intensity and texture detail information. Complementary information measurement is applied in the loss function to improve the preservation of fusion images’ thermal radiation and detail characteristics.
@@
    A dual attention fusion block is designed for information integration based on spatial and channel attention mechanisms, where the vital spatial region and channel importance are determined by the attention map for multi-scale feature fusion.
@@
    Complementary information of infrared and visible images is effectively measured through the statistical methods and generates adaptive weight coefficients for each item in the loss function to constrain the difference of fusion results and input image pairs, quantifying the degree of the vital information retention of each modality.
@@
    Extensive experiments are implemented on the TNO and RoadScene datasets, illustrating that our complementary information measurement applied multi-scale network obtains competitive performances compared with seven state-of-the-art infrared and visible image fusion methods in terms of qualitative and quantitative comparisons. Ablation studies show the effectiveness of our information measurement and integration method.
@@
The remainder of this article is organized as follows. Section II introduced previous related works. In Section III, we propose multi-scale dual attention (MDA) framework for infrared and visible image fusion. Extensive experiments compared with seven methods on the TNO and RoadScene datasets are illustrated in Section IV. Finally, Section V shows the conclusion.
II Related Work
II-1 Infrared and visible image fusion methods
@@
With the rapid development of deep learning, end-to-end image fusion methods applying neural networks have emerged, which can be divided into AE-based, CNN-based and GAN-based methods. Early AE-based methods [21, 12] extracted features and reconstructed images through an encoder and decoder pretrained on the MS-COCO dataset, and integrated features via traditional fusion methods like addition, max selection and L1-norm fusion strategies. However, downsample operation in AE-based methods results in losing detailed information, meanwhile training on the MS-COCO dataset would make it difficult for the model to learn the property of infrared images. [22, 23] add short connections between the encoder and decoder to preserve detailed information. Part of the CNN-based methods [24] adopted addition, weighted average and L1-norm to fuse features, while the others [13, 25] concatenate features on the channel dimension to aggregate information. These kinds of methods fulfill feature extraction, fusion and reconstruction end-to-end without any downsample operations, and the fusion module learns to integrate modal-wise information through a well-designed loss function. FusionGAN [11] firstly established an adversarial game between a generator and a discriminator to keep the thermal intensity information and visible detail information simultaneously, implementing feature extraction, fusion and reconstruction via the implicit way. However, the balance between the generator and a single discriminator is hard to balance during the training. DDcGAN [26] and SDDGAN [27] adopted a generator to fuse the image and a dual-discriminator to force the fused image to retain intensity and detail information separately. AttentionFGAN [28] and MFEIF[29] introduce the attention mechanism into the generator and discriminator to force the fusion results focusing on the most discriminator target regions. YDTR[30] extracted local features and significant context information via a Y-shape dynamic Transformer. Indeed, the multi-task image fusion framework has also been proposed, such as UMF-CMGR[31] combines cross-modality image alignment and fusion to avoid suffering ghosts of misaligned cross-modality image pairs.
II-2 Complementary information measurement
@@ III Methods
III-A Network Architecture
In this section, we present the network architecture of the proposed framework, as shown in Fig. 1. As can be seen, our method contains three main parts: 1. Residual downsample block applies multi-resolution residual convolution streams for extracting both detail-rich and semantic-rich feature representations, 2. The spatial and channel attention-based dual attention fusion block aggregates feature from both the same and adjacent scales, 3. Residual reconstruction block for image reconstruction by using the multi-scale fused features. The network as a whole is a two-branch structure, extracting multi-scale features of infrared and visible images respectively.
@@ III-A Network Architecture
III-B Dual Attention Fusion Block
@@
where $MLP$ denotes two fully connected layers, and $\sigma$ denotes sigmoid function. The above operation enables the fusion features adaptively select the inter-channel complementary information.
@@ III-C Loss Function
$I_{f}^{W}$, $I_{ir}^{W}$ and $I_{vis}^{W}$ denote the fusion image, infrared image and visible image in the local sliding window $W$, respectively. Considering the area of targets with high temperatures, we set the size of the sliding window to 21 ×21 in this study. Pixel intensity in a small area is mainly dominated by the gray value rather than entropy, thus we replace entropy with the mean pixel value in the window, and the weight calculation of patch loss is as follows:
@@
where $\gamma$ controls the trade-off. Except for the constraint between input and fusion result, we also set hierarchical loss for each fusion block to guarantee the training efficiency of the dual attention module. Therefore, $L_{feature}$ is defined as follows:
@@ III-D1 Data
In the training phase, we use KAIST [53] dataset to train our network, which consists of nearly 95k aligned IR-VIS image pairs (640×480) taken from the vehicle. We crop source images to patches of size 192 × 192 and transfer the color space of visible images to YCbCr. Since the infrared images are grayscale, we only use the Y channel to fuse with infrared images. In the testing phase, TNO  [^1] https://figshare.com/articles/TNO_Image_Fusion_Dataset/1008029. dataset [54] and RoadScene  [^2] https://github.com/hanna-xu/RoadScene. dataset provided by [34] is used to evaluate the performance of our method. TNO is a classical IR-VIS dataset with rich scenes, and we choose 51 image pairs to validate the generalization of our method. RoadScene contains 221 aligned IR and VIS image pairs, which are highly representative scenes from the FLIR video, and those image pairs are aligned accurately while denoised.
III-D2 Training details
@@
The batch size and epoch are set to 24 and 5, respectively. We set $\alpha=1e-8$, $\beta=1e7$, $\gamma=2$ and $\eta=0.02$ in the loss functions, respectively. For the complementary information measurement parameters in $L_{image}$, the $\delta$, $\zeta$ and $c$ is set as $0.167$, $20$ and $3\times 10^{3}$, respectively. We optimize the loss function using the Adam with a learning rate $1\times 10^{-5}$. Experiments are performed on NVIDIA GeForce RTX 3090 GPU and 2.90GHz Intel Xeon Gold 6226R CPU.
Figure 4: Qualitative comparison of our method with 7 state-of-the-art models on six infrared and visible image pairs in the TNO dataset. The first and second rows are visible and infrared images, respectively. From the third to the last rows are fusion results of DenseFuse[12], DRF[38], RFN-Nest[23], GANMcC[15], STDFusionNet[40], UMF-CMGR[31], YDTR[30] and our MDA.
IV Experiments

### v2

@@ Abstract
Infrared and visible image fusion aims at generating a fused image containing the intensity and detail information of source images, and the key issue is effectively measuring and integrating the complementary information of multi-modality images from the same scene. Existing methods mostly adopt a simple weight in the loss function to decide the information retention of each modality rather than adaptively measuring complementary information for different image pairs. In this study, we propose a multi-scale dual attention (MDA) framework for infrared and visible image fusion, which is designed to measure and integrate multi-scale complementary information in both structure and loss function at the image and patch level. In our method, the residual downsample block decomposes source images into three scales first. Then, dual attention fusion block integrates complementary information and generates a spatial and channel attention map at the same and adjacent scale for feature fusion. Finally, the output image is reconstructed by the residual reconstruction block. Loss function consists of image-level, feature-level and patch-level three parts, of which the calculation of the image-level and patch-level two parts are based on the weights generated by the complementary information measurement. Indeed, to constrain the pixel intensity distribution between the output and infrared image, a style loss is added. Our fusion results perform robust and informative across different scenarios. Qualitative and quantitative results on three datasets illustrate that our method is able to preserve both thermal radiation and detailed information from two modalities and achieve comparable results compared with the other state-of-the-art methods. Ablation experiments show the effectiveness of our information integration architecture and adaptively measure complementary information retention in the loss function. Our code is available at https://github.com/SSyangguang/MDA.
keywords:
@@
Infrared and visible image fusion , Dual attention , Adaptive weight generation
journal: Neurocomputing
[…]
[1]organization=State Key Laboratory of Integrated Services Networks, School of Electronic Engineering, Xidian University, city=Xi’an, postcode=Shaanxi 710071, country=China
[…]
[2]organization=Chongqing Key Laboratory of Image Cognition, Chongqing University of Posts and Telecommunications, city=Chongqing, postcode=Chongqing 400065, country=China
1 Introduction
@@
Due to the limitations of different physical sensors, images from one single sensor is hard to comprehensively describe the scene information. Image fusion is one kind of multimodal information fusion technique, which aims at integrating images from different modalities into a single more informative image at the pixel level [1, 2, 3]. Infrared and visible image fusion is a branch of the image fusion techniques to generate an image containing both thermal and detail information, in which the infrared images can easily distinguish targets with high temperature from the background [4]. However, infrared images typically have low spatial resolution and poor details. As complementary information to the infrared images, the visible images contain rich textures but are sensitive to unfavorable illumination conditions like nighttime and bad weather. The fusion results of infrared and visible images can depict the scene from aspects of thermal radiations and the human visual system simultaneously, which is practical in many fields, such as recognition, surveillance and object detection [5, 6, 7].
@@
According to the adopted theories [8, 9, 10], the traditional image fusion method can be divided into several categories, namely multi-scale transform-, sparse representation-, subspace-, saliency-bases methods, hybrid models, and other models. However, the fusion performance of these methods is highly related to their handcrafted features and fusion schemes [11]. In recent years, deep learning-based methods have been successfully applied to image fusion with their strong feature extraction ability and adaptive fusion strategy, which overcome the drawbacks of traditional methods to a certain extent. The overall deep learning-based image fusion pipeline can be processed implicitly by an end-to-end convolutional neural network under the well-designed network architecture and loss function, which is able to characterize complex relationships between input image pairs and the fusion image[12]. As suggested by [13], the deep learning-based image fusion methods can be divided into autoencoder (AE)-based methods, convolutional neural network (CNN)-based methods and generative adversarial network (GAN)-based methods. Among them, GAN-based methods rely on an adversarial model between the generator and discriminator to implicitly fulfill feature extraction, feature fusion and image reconstruction [14, 15]. The AE-based and CNN-based methods generally use an autoencoder or fully convolutional network to realize the feature extraction and reconstruction, while the fusion strategy is conventional fusion rules or concatenation on channel dimension, such as DenseFuse [16] and PMGI [17].
@@
Infrared and visible images capture radiation and illumination information respectively, and typically pixel intensity is used to constrain thermal radiation information retention while gradient variation is used to constrain texture detail information [18]. Nevertheless, the complementary information varies greatly in different scenes. For instance, some infrared images achieve better detail at night, which makes the above information measurement method not accurate enough [19]. Existing methods prefer preserving pixel intensity and detail information via the mean square error and structural similarity in the loss function separately, which did not consider their information retention-based adaptive weight coefficients before terms in the loss function. Most of them simply add residual or dense connection in the network architecture to capture texture details, failing to measure the multi-scale complementary information among different modalities and scales, which is a key issue in image fusion [20]. Additionally, no ground truth in the infrared and visible image fusion task makes manually designed information retention methods unreasonable. Therefore, determining the degree of vital information retention based on each modality’s own attributes and effectively integrating complementary information makes infrared and visible image fusion more reasonable. The existing methods barely applied complementary information measurement neither in calculating loss function nor feature fusion module, and the cooperation of the multi-scale complementary information between loss function and fusion structure is insufficient. U2Fusion [21] adopted complementary information measurement in the loss function, but it did not consider measurement of the multi-scale information, such as the image-level and patch-level information capturing global and local complementary information.
@@
[22] showed that multi-scale analysis is similar to the human visual characteristics, and the fused image with such property is capable of capturing contextual information from the high-level representations and detailed information from the low-level representations [23]. In this paper, we propose a multi-scale dual attention (MDA) framework applying multi-scale complementary information both in loss function and network architecture to effectively measure and integrate vital information at three scales. The given images are first decomposed into three scales via the residual downsample block. High-resolution features are more beneficial for measuring and preserving detailed information, while the pixel intensity distribution can be measured accurately at low resolution. Differing from those directly aggregate features by addition or concatenation, we introduce channel and spatial attention mechanisms to exploit ’what’ and ’where’ is informative for the given infrared-visible image pairs [24, 25], and then fuse features extracted by the residual downsample block based on the attention maps. Dual attention fusion block merges features from the same and adjacent scale of different modalities along the spatial and channel dimensions, each of them generating an attention map for feature aggregation, which is considered the multi-scale complementary information interaction process. Meanwhile, each of the fusion blocks is constrained by the hierarchical loss to ensure the efficiency of the complementary information interaction and integration between the same and different scales. Fused features are upsampled to the same size as the input image and reconstructed by the residual reconstruction block.
@@
As an auxiliary way to fuse the multi-scale complementary information in the network structure, the pixel item in the loss function consists of intensity and detail two parts with different adaptive weight coefficients obtained by the complementary information measurement. The generated coefficients are calculated based on the full resolution inputs and local patch, which is considered as the multi-scale complementary information assistance in the loss function. For the semantic-rich representations, we adopt statistical methods to generate weight coefficients for the pixel intensity term in the loss function. For the detail-rich representations, we apply the gradient operator on the input image pairs to measure the retention of the texture information, and then weight coefficients of the detail term are generated. The above weight generation method is applied to the image-level and patch-level simultaneously to assist multi-scale information in the network structure, enabling the multi-scale complementary information interaction efficiently. Additionally, we add a style loss to constrain pixel intensity distribution differences between the fusion image and the infrared image. The main contributions in this work are as follows:
@@
    We design a multi-scale dual attention (MDA) framework for infrared and visible image fusion that measures and fuses multi-scale complementary information from different modalities across multiple spatial scales, both utilizing pixel intensity and texture detail information. The network decomposes the image pairs into three scales to capture semantic-rich and detail-rich information, meanwhile the image-level and patch-level constraints are applied to assist the multi-scale complementary information integration.
@@
    A dual attention fusion block is designed for information integration based on spatial and channel attention mechanisms, where the vital spatial region and channel importance are determined by the attention map for multi-scale feature fusion. Feature maps of the same and adjacent scales are fused via the dual attention fusion block to ensure the efficiency of the multi-scale information interaction.
@@
    Complementary information of image-level and patch-level infrared and visible representations is effectively measured through the statistical methods and gradient operator and generates adaptive weight coefficients for each item in the loss function to constrain the difference of fusion results and input image pairs, quantifying the degree of the vital information retention of thermal radiation and detail characteristics.
@@
    Extensive experiments are implemented on the TNO, RoadScene and $M^{3}FD$ datasets, illustrating that our complementary information measurement applied multi-scale network obtains competitive performances compared with seven state-of-the-art infrared and visible image fusion methods in terms of qualitative and quantitative comparisons. Ablation studies show the effectiveness of our multi-scale information measurement and integration method.
@@
The remainder of this article is organized as follows. Section II introduced previous related works. In Section III, we propose a multi-scale dual attention (MDA) framework for infrared and visible image fusion. Extensive experiments compared with seven methods on the TNO, RoadScene and $M^{3}FD$ datasets are illustrated in Section IV. Finally, Section V shows the conclusion.
2 Related Work
2.1 Infrared and visible image fusion methods
@@
With the rapid development of deep learning, end-to-end image fusion methods applying neural networks have emerged, which can be divided into AE-based, CNN-based and GAN-based methods. Early AE-based methods [26, 16] extracted features and reconstructed images through an encoder and decoder pretrained on the MS-COCO dataset, and integrated features via traditional fusion methods like addition, max selection and L1-norm fusion strategies. However, downsample operation in AE-based methods results in losing detailed information, meanwhile training on the MS-COCO dataset would make it difficult for the model to learn the property of infrared images. [27, 28] add short connections between the encoder and decoder to preserve detailed information. Part of the CNN-based methods [29] adopted addition, weighted average and L1-norm to fuse features, while the others [17, 30] concatenate features on the channel dimension to aggregate information. These kinds of methods fulfill feature extraction, fusion and reconstruction end-to-end without any downsample operations, and the fusion module learns to integrate modal-wise information through a well-designed loss function. FusionGAN [14] firstly established an adversarial game between a generator and a discriminator to keep the thermal intensity information and visible detail information simultaneously, implementing feature extraction, fusion and reconstruction via the implicit way. However, the balance between the generator and a single discriminator is hard to balance during the training. DDcGAN [31] and SDDGAN [32] adopted a generator to fuse the image and a dual-discriminator to force the fused image to retain intensity and detail information separately. AttentionFGAN [33] and MFEIF[34] introduce the attention mechanism into the generator and discriminator to force the fusion results focusing on the most discriminator target regions. YDTR[35] extracted local features and significant context information via a Y-shape dynamic Transformer. SwinFusion [36] devised a general image fusion framework based on the cross-domain long-range learning and Transformer. LRRNet [37] introduced learnable image decomposition to guide the construction of the network architecture. FAFusion[38] leveraged low- and high-frequency information for preserving the global brightness and local texture details of the infrared and visible images. Indeed, the multi-task image fusion framework has also been proposed, such as UMF-CMGR[39] combined cross-modality image alignment and fusion to avoid suffering ghosts of misaligned cross-modality image pairs. SeAFusion [40] and PSFusion [41] combined image fusion tasks with high-level vision tasks to introduce semantic information from the semantic segmentation, which can improve the performance of both tasks simultaneously.
2.2 Complementary information measurement
@@ 3 Methods
3.1 Network Architecture
In this section, we present the network architecture of the proposed framework, as shown in Fig. 1. As can be seen, our method contains three main parts: 1. Residual downsample block applies multi-resolution residual convolution streams for extracting both detail-rich and semantic-rich feature representations, 2. The spatial and channel attention-based dual attention fusion block aggregates features from both the same and adjacent scales, 3. Residual reconstruction block for image reconstruction by using the multi-scale fused features. The network as a whole is a two-branch structure, extracting multi-scale features of infrared and visible images respectively.
@@ 3.1 Network Architecture
3.2 Dual Attention Fusion Block
@@
where $MLP$ denotes two fully connected layers, and $\sigma$ denotes sigmoid function. The above operation enables the fusion features to adaptively select the inter-channel complementary information.
@@ 3.3 Loss Function
$I_{f}^{W}$, $I_{ir}^{W}$ and $I_{vis}^{W}$ denote the fusion image, infrared image and visible image in the local sliding window $W$, respectively. Considering the area of targets with high temperatures existing at boundaries between patches, we set the size of the sliding window to 21 ×21 and the overlap pixel between the sliding window is set to 1 in this study. Pixel intensity in a small area is mainly dominated by the gray value rather than entropy, thus we replace entropy with the mean pixel value in the window, and the weight calculation of patch loss is as follows:
@@
where $\gamma$ controls the trade-off. Except for the constraint between input and fusion result, we also set hierarchical loss for each fusion block to guarantee the training efficiency of the dual attention module and multi-scale information interaction. Therefore, $L_{feature}$ is defined as follows:
@@ 3.4.1 Data
In the training phase, we use KAIST [62] dataset to train our network, which consists of nearly 95k aligned IR-VIS image pairs (640×480) taken from the vehicle. We crop source images to patches of size 192 × 192 and transfer the color space of visible images to YCbCr. Since the infrared images are grayscale, we only use the Y channel to fuse with infrared images. In the testing phase, TNO  [^1] https://figshare.com/articles/TNO_Image_Fusion_Dataset/1008029. dataset [63], RoadScene [^2] https://github.com/hanna-xu/RoadScene. dataset [44] and $M^{3}FD$ [^3] https://github.com/JinyuanLiu-CV/TarDAL. dataset provided by [64] is used to evaluate the performance of our method. TNO is a classical IR-VIS dataset with rich scenes. RoadScene contains 221 aligned IR and VIS image pairs, which are highly representative scenes from the FLIR video, and those image pairs are aligned accurately while denoised. $M^{3}FD$ is a high-resolution IR-VIS dataset consisting of 4200 aligned pairs with six classes of annotated objects, and 300 pairs are chosen to evaluate the performance of methods.
3.4.2 Training details
@@
The batch size and epoch are set to 24 and 5, respectively. We set $\alpha=1e-8$, $\beta=1e5$, $\gamma=2$ and $\eta=0.02$ in the loss functions, respectively. For the complementary information measurement parameters in $L_{image}$, the $\delta$, $\zeta$ and $c$ is set as $0.167$, $20$ and $3\times 10^{3}$, respectively. We optimize the loss function using the Adam with a learning rate $1\times 10^{-5}$. Experiments are performed on NVIDIA GeForce RTX 3090 GPU and 2.90GHz Intel Xeon Gold 6226R CPU.
4 Experiments
4.1 Implementation